# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v1

@@ 8 Discussion
our exploration of TRiSM-based governance for LLM-powered Agentic AI systems reveals critical insights into technical design, ethical oversight, regulatory alignment, and future challenges. Below, we discuss the broader implications of our findings, structured into key areas for clarity.
@@ 8.1 Technical Implications of TRiSM for LLM Agent Design
The AI TRiSM framework (Trust, Risk, and Security Management) imposes concrete technical requirements on how autonomous LLM-driven agents are built and deployed. A core implication is the need to embed real-time monitoring and control mechanisms into agent architectures. Rather than treating LLMs agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” guardrails_ai_guardrails_2024 . For example, there is discusison on designing specialized “guardian agents” within AMAS securiti2025aitrism . In this paradigm, such agents serve as proactive monitors that filter sensitive data and establish baselines of normal behavior, while operator agents dynamically enforce policies at runtime (e.g. blocking disallowed actions such as outputting personally identifiable information). This layered control strategy transforms the technical architecture: an autonomous LLM agent is now supplemented by meta-agents that supervise its inputs, outputs, and tool use in real time.
Prior research highlights risks like “excessive agency” owasp2025agentic where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals). TRiSM-driven agentic design mitigates these failure modes by constraining agent autonomy within well-defined safety bounds. Likewise, emerging threats specific to Agentic AI such as prompt injection kosinski2024prompt attacks, memory poisoning atkins2023memorypoisoning , or cascading hallucinations huang_survey_2023 underscore the need for built-in risk controls. By incorporating anomaly detection and policy-checking modules (as per the Sentinel/Operator model), an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g. masking a sensitive datum or stopping an unsafe action). This aligns with calls in the security community for proactive measures in Agentic AI, blending traditional cybersecurity techniques (access control, logging, sandboxing) with AI-specific safeguards like LLM firewalls and adversarial robustness checks raza-etal-2022-accuracy .
In summary, TRiSM’s technical implications mean that autonomous LLM agents should no longer be deployed as stand-alone intelligent actors; instead, they operate under an active governance fabric of monitors, validators, and enforcement agents that ensure trustworthiness and safety by design.
@@ 8.2 Ethical and Societal Ramifications of Multi-Agent AI
Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions. Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness raza2025responsible all of which are vital for public trust in AI systems. A central concern is accountability: when AI agents make autonomous decisions that affect humans, who is answerable for the outcomes? TRiSM-based governance insists that organizations retain clear responsibility for their AI’s actions, rather than obscuring blame behind algorithmic “black boxes”. This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed gu2024artificial . Recent guidance on trustworthy AI chen2024trustworthy often highlights accountability and explainability as key pillars of trust.
In practice, our approach means each autonomous agent’s decisions should be transparent enough to be understood and challenged by human reviewers when necessary. Human oversight is another ethical imperative tightly coupled with accountability. TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under defined governance. Human operators or “AI managers” must have the ability to intervene or override when an agent’s behavior deviates from acceptable bounds or when moral judgment is required. Indeed, high-level policy frameworks (such as the EU’s AI ethics guidelines euregulationai2024 ) explicitly call for “human agency and oversight” in AI systems.
In multi-agent setups, this may involve dashboard interfaces where humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly feng2025multi . The risk of “user complacency”, trusting an autonomous agent too much , has been noted as a hazard parasuraman2010complacency . TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision. Fairness and bias mitigation are also critical societal considerations.
Our governance approach therefore incorporates bias audits and fairness checks throughout the agent lifecycle raza_fair_2024 . Techniques like pre-deployment bias testing, continuous monitoring for disparate impacts, and diverse stakeholder evaluation panels can be employed. These measures echo regulatory expectations; the EU’s AI Act euregulationai2024 and related guidelines enumerate “diversity, non-discrimination and fairness” as core requirements for trustworthy AI. In deploying LLM-based agents, we must ensure they do not treat individuals or groups inequitably, for example, content-filtering agents should apply policies uniformly across demographic groups, and task-planning agents should not propagate historical biases in resource allocation decisions.
In sum, TRiSM-oriented governance extends beyond preventing technical failures: it seeks to uphold ethical norms and human rights, ensuring that autonomy in AI does not come at the expense of justice, transparency, or human dignity.
@@ 8.3 Alignment with Emerging AI Regulations and Standards
The principles embedded in TRiSM align closely with emerging regulatory frameworks for AI. This convergence means that adopting TRiSM-based governance can help organizations meet new legal obligations and industry standards. For example, the European Union’s AI Act eu_ai_act (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems. These are precisely the capabilities that a TRiSM approach cultivates.
By instituting continuous risk assessment, documentation of AI decision processes, and oversight mechanisms, a TRiSM-governed multi-agent system inherently addresses many of the EU Act’s requirements (e.g. having a risk management system and post-market monitoring for AI. Notably, the Act also stresses accuracy, robustness, and cybersecurity for high-risk AI, qualities that TRiSM’s security management component is designed to ensure (through adversarial resilience, access control, etc.). Similarly, international AI governance standards are emerging that mirror TRiSM’s tenets. ISO/IEC 42001:2023, the first global standard for AI management systems, highlights requirements such as transparency, accountability, bias mitigation, safety, and privacy in AI development
TRiSM’s trust and risk management focus naturally encompasses these elements: for instance, trust in TRiSM relates to reliable, truthful outputs (promoting transparency), while explicit risk management aligns with accountability for negative outcomes. By implementing TRiSM, organizations essentially put in place the processes that ISO 42001 iso42001 and similar standards call for (e.g. leadership oversight, documented risk controls, ongoing monitoring and improvement cycles). Another example is the U.S. NIST AI Risk Management Framework dotan2024evolving , which emphasizes many of the same concepts: identifying risks, embedding governance, and cultivating trustworthiness in AI.
By following TRiSM guidelines: for example, maintaining an “AI catalog” of all models/agents in use and their purposes, enforcing policies via sentinel/operator agents, and logging every AI decision, organizations create an audit-ready environment. In the event of an incident or an inquiry by authorities, they can demonstrate traceability and control over their autonomous agents, which will be crucial for regulatory compliance and liability management.
@@ 8.4 Limitations and Current Research Gaps
While the TRiSM-based approach appears promising, our work also revealed several limitations and open challenges in current research.
First, limited benchmark evaluations pose a problem. The AI safety and agent governance community lacks widely accepted benchmarks to quantitatively assess trustworthiness or risk in multi-agent LLM systems. Unlike classical AI domains (vision, NLP) that have standard test suites, there is no consensus on how to measure an “AI agent” ability to operate safely under TRiSM principles. This makes it difficult to compare different governance strategies or to track progress objectively. We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations standen2025adversarial or “red-teaming” exercises ganguli_red_2022 that can stress-test agentic systems and score their resilience (e.g. measuring success rates of prompt injection attacks or frequency of policy violations caught by oversight agents).
Secondly, there is a paucity of real-world validation for many TRiSM-inspired controls. Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks charli_trism_2024 . It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments. Additionally, integrating TRiSM with legacy systems poses practical challenges , for example, the earlier work notes compatibility issues when bolting on trust/security layers to existing AI pipelines. This suggests a limitation in how easily current AI deployments can fit TRiSM controls, a topic that merits further engineering research.
Another critical gap is adversarial robustness. As we improve defenses, attackers will inevitably adapt. Recent findings show that LLM-based systems remain vulnerable to cleverly crafted attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails raza2025justhumansneedvaccines . TRiSM solutions today are not foolproof against these. For example, an agent designed to mask secret data might itself be tricked into revealing it if the oversight logic fails to anticipate a new attack pattern. The literature identifies “evolving threats” and “adversarial attacks” as ongoing obstacles to trustworthy AI securiti2025aitrism . This underlines a need for continuous updates and adaptive security in any TRiSM implementation.
@@ 8.5 Future Roadmap for Agentic AI TRiSM
Future Roadmap for Agentic AI Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Benchmarking & Evaluation Cognitive Capability Expansion
Figure 3: Strategic roadmap for LLM-enabled Agentic AI, grouped into eight priority research and development domains.
Drawing on our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure 3. These recommendations span both technical system design improvements and governance-level policy initiatives:
  • Develop Standardized Evaluation Benchmarks: The community should create open benchmarks and challenge environments to test multi-agent AI governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed agent system performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
  • Advance Adversarial Robustness Techniques: Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” bacudio2011overview , and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames schneider2020wargaming , can help discover vulnerabilities in a controlled way before real adversaries do.
  • Human-Centered Oversight Tools: We encourage designing better interfaces and protocols for human oversight of Agentic AI. Borrowing from human-computer interaction AmershiWVFNCSIB19 and cognitive engineering hexmoor2025behaviour , researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
  • Regulatory Sandboxes and Compliance-by-Design: Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations. Moreover, adopting a compliance-by-design mindset is key: future AI system designs should bake in the requirements of frameworks like the EU AI Act and ISO 42001 from the startrather than retrofit them.
@@ 9 Conclusion
TRiSM-based governance offers a promising scaffold to ensure that autonomous LLM-powered agents are trustworthy, accountable, and secure. Our discussion has analyzed how this framework influences technical design decisions, mandates ethical guardrails, and dovetails with emerging regulatory regimes. While current research is nascent and not without limitations, the path forward is clear. By rigorously testing these systems, strengthening them against adversaries, and crafting policies and standards in tandem with technological advances, we can enable powerful multi-agent AI systems to operate beneficially under robust oversight. The stakes are high but with a proactive, interdisciplinary approach, we can achieve a balance where innovation in AI goes hand-in-hand with responsibility and trust. As future work tackles the open challenges identified, we anticipate that TRiSM principles will transition from conceptual best-practice to standard operating procedure for Agentic AI, ensuring these systems earn and maintain the confidence of all stakeholders involved.

### v2

@@ 9 Discussion
We discuss the broader implications of our findings below:
@@ 9.1 Technical Implications of TRiSM for LLM Agent Design
Rather than treating LLM agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” [216]. For example, there is discussion on designing specialized “guardian agents” within AMAS [40]. In this paradigm, such agents serve as proactive monitors that filter sensitive data and establish baselines of normal behavior, while operator agents dynamically enforce policies at runtime (e.g. blocking disallowed actions such as outputting personally identifiable information).
Prior research highlights risks like “excessive agency” [106] where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals). TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds. Likewise, emerging threats specific to Agentic AI systems such as prompt injection [23] attacks, memory poisoning [137], or cascading hallucinations [217] highlights the need for more built in risk controls. By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g. masking a sensitive datum or stopping an unsafe action) .
@@ 9.2 Ethical and Societal Ramifications of Multi-Agent AI
Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions. Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness [169], all of which are vital for public trust in AI systems. Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”. This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed [161]. Recent guidance on trustworthy AI [172] often highlights accountability and explainability as key pillars of trust.
In practice, this review sheds some light on how each autonomous agent decisions should be transparent enough to be understood and challenged by human reviewers when necessary [218]. Human oversight is another ethical imperative tightly coupled with accountability. TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under a defined governance framework [219].
In MAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly [126]. The risk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [220]. TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision.
@@ 9.3 Alignment with Emerging AI Regulations and Standards
The principles embedded in TRiSM align closely with emerging regulatory frameworks for AI. This convergence means that adopting TRiSM-based governance can help organizations meet new legal obligations and industry standards. For example, the European Union’s AI Act [94] (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems. These are precisely the capabilities that a TRiSM approach promotes.
The EU AI Act also stresses accuracy, robustness, and cybersecurity for high-risk AI, qualities that TRiSM security management component is designed to ensure (through adversarial resilience, access control, etc.). Similarly, international AI governance standards are emerging that mirror TRiSM’s principles. ISO/IEC 42001:2023, the first global standard for AI management systems, highlights requirements such as transparency, accountability, bias mitigation, safety, and privacy in AI development. By implementing TRiSM, organizations essentially put in place the processes that ISO 42001 [89] and similar standards call for (e.g. leadership oversight, documented risk controls, ongoing monitoring and improvement cycles). Another example is the U.S. NIST AI Risk Management Framework [87], which emphasizes many of the same concepts: identifying risks, embedding governance, and cultivating trustworthiness in AI.
@@ 9.4 Limitations and Current Research Gaps
While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research. First, the classical AI domains (vision, NLP) that have standard test suites, there is no consensus on how to measure an “AI agent” ability to operate safely under TRiSM principles. This makes it difficult to compare different governance strategies or to track progress objectively. We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [210] or “red-teaming” exercises [221] that can stress-test Agentic AI systems and score their resilience.
Secondly, there is a lack of real-world validation for many TRiSM-inspired controls. Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [31]. It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments.
Another critical gap is adversarial robustness. As we improve defenses, attackers will inevitably adapt [222]. Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [25]. TRiSM solutions today are not foolproof against these. We identify evolving threats and adversarial attacks as ongoing obstacles to trustworthy AI [40].
@@ 9.5 Future Roadmap for Agentic AI TRiSM
Table 9: Future Roadmap for Agentic AI: Eight Strategic Domains
Scalable Agent Architectures Secure Multi-Agent Collaboration
Explainable Multi-Agent Decisions Human-Centered Trust Design
Autonomous Lifecycle Management Ethics & Governance
Benchmarking & Evaluation Cognitive Capability Expansion
Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure LABEL:fig:roadmap_mindmap. These recommendations span both technical system design improvements and governance-level policy initiatives:
Develop Standardized Evaluation Benchmarks. The community should create open benchmarks and challenge environments to test multi-agent AI governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed agent system performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
Advance Adversarial Robustness Techniques. Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” [223], and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [224] , can help discover vulnerabilities in a controlled way before real adversaries do.
Human-Centered Oversight Tools. We encourage designing better interfaces and protocols for human oversight of Agentic AI systems. Borrowing from human-computer interaction [225] and cognitive engineering [226], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
Regulatory Sandboxes and Compliance-by-Design: Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations. Moreover, adopting a compliance-by-design mindset is key: future AI system designs should bake in the requirements of frameworks like the EU AI Act and ISO 42001 from the start rather than retrofit them.
@@ 10 Conclusion
The rapid advancement of Agentic AI, driven by LLM-based multi-agent systems, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services. However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations. This review has systematically explored these challenges, proposing a TRiSM framework structured around four pillars: Governance, Explainability, ModelOps, and Privacy/Security, specifically adapted for LLM-based Agentic Multi-Agent Systems (AMAS).
Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency. The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities.
Looking forward, the proposed research roadmap emphasizes scalable, verifiable, and regulation-compliant Agentic AI systems. Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination. By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix
Table A.1: Key terminologies for LLM-based agentic AI systems.
Term Definition
Agentic AI System A multi-agent architecture powered by large language models (LLMs), where autonomous agents collaborate, plan, and execute tasks over extended horizons with persistent memory and dynamic role assignment.
Autonomy Model The mechanism by which an agent decides and acts without direct human intervention, often using goal-driven planning and chain-of-thought reasoning.
Chain of Thought (CoT) A prompting strategy in which an LLM generates intermediate reasoning steps before producing a final answer or action, enhancing interpretability and multi-step planning.
Counterfactual Analysis An interpretability technique that examines how altering certain inputs or agent contributions would change the overall system outcome, revealing causal dependencies among agents.
Explainability The capacity of an AI system (or individual agent) to produce human-understandable justifications or rationales for its decisions and actions, often via LIME, SHAP, or decision provenance graphs.
Foundation Model (LLM) A pretrained large language model (e.g., GPT-4, LLaMA) that serves as the “brain” of each agent, providing generative capabilities, reasoning, and tool-calling support.
Shared Memory (Persistent Memory) A centralized or distributed store (often a vector database) where agents write and retrieve contextual information, enabling long-term planning and consistency across iterations.
ModelOps The practice of managing AI models (and agent prompts) throughout their lifecycle—development, deployment, monitoring, and retirement—with version control, CI/CD testing, and drift detection.
Application Security Safeguards and best practices (e.g., prompt sanitation, authentication, sandboxing) designed to protect agentic systems from prompt injection, identity spoofing, and lateral exploits.
Model Privacy Techniques (e.g., differential privacy, homomorphic encryption, secure enclaves) that ensure sensitive data—either during training or inter-agent communication—remains protected in multi-agent workflows.
Prompt Injection A security exploit in which an attacker crafts input containing hidden instructions that corrupt an agent’s reasoning or propagate malicious commands through agent interactions (“prompt infection”).
Retrieval-Augmented Generation (RAG) A framework where agents query an external knowledge store (e.g., vector database) to fetch relevant documents or facts, then condition their LLM responses on that retrieved context.
Role-Specialized Coordination An architectural pattern in which each agent is assigned a specific function (e.g., planner, verifier, coder) and collaborates via structured communication protocols to achieve complex tasks.
Decision Provenance Graph A graph-based representation that traces data flows and decision steps across multiple agents, enabling post-hoc auditing and system-level interpretability.
Tool-Use Interface The mechanism by which an agent issues structured commands (e.g., API calls, code execution) to external services or environments and incorporates the results back into its reasoning.
Trust Score A composite metric that quantifies an agent’s reliability, alignment with user goals, and consistency over time, often combining accuracy, safety-violation rates, and calibration of confidence.
Composite Metric An aggregate evaluation score (e.g., a weighted sum of trustworthiness, explainability, user-centered performance, and coordination metrics) used to benchmark different agentic systems.
Figure A.1: Traditional AI Agent vs. Agentic AI.